b1az3:

The data list shows the cost, in dollars, of different brands of sneakers at a shoe store. 44.99, 59.99, 32.30, 65.00, 52.59 What is the population standard deviation, to the nearest tenth, of the data (assuming this is all of the data)?

AZ:

All the progress we've made so far was on here: https://questioncove.com/study#/updates/6048a70c3ad7c95181004c00 \(\sigma =\sqrt{ \dfrac{\sum(x_i-\mu)^2}{N}}\) where \(\sigma\) = population standard deviation \(\sum\) = the sum from 1 to N \( x_i\) = each individual data point \(\mu\) = mean of the data list N = total number of data points

AZ:

You eventually after much labor and direction found that the mean \(\mu = 50.974\) so now we need to calculate \( (x_i - \mu)\) and then we can find \( (x_i - \mu)^2\) so where we left off, what is: 44.99 - 50.974 = 59.99 - 50.974 = 32.30 - 50.974 = 65.00 - 50.974 = 52.59 - 50.974 = tell me the numbers for each subtraction